OpenStudy (anonymous):

Given f(x) = x^2+2, find f(a+2) − f(a)

OpenStudy (anonymous):

4a+4

OpenStudy (anonymous):

f(a+2)=(a+2)^2+2 f(a)=a^2+2 f(a+2)-f(a)=(a+2)^2+2-(a^2+2)

OpenStudy (anonymous):

Which one is correct? :(

OpenStudy (anonymous):

brusmack just shows how to do it. same answer

OpenStudy (anonymous):

THANK YOUUUUUUUUUUUUUUUUUUUUU!